OpenStudy (anonymous):

Sasha is going to choose 4 kinds of fruit out of the 9 kinds available at the grocery store. In how many different ways can she do that?

OpenStudy (anonymous):

9*8*7*6 = 3024

OpenStudy (shamim):

do u hv idea abt combination

OpenStudy (epoweritheta):

9C4 = 9.8.7.6/1.2.3.4=126

OpenStudy (shamim):

nCr=n!/(r!*(n-r)!)